Comerpickles:

The state of New Jersey has a population of about 8.9 million and its area is about 8700 mi². What is the population density of New Jersey? Round your answer to the nearest whole number. Question 1 options: 1,023 102 890 100

Comerpickles:

@563blackghost

Comerpickles:

@Vocaloid

563blackghost:

\(\large\bf{Population~Density = \frac{number~of~people}{area}}\)

Comerpickles:

977.5280..

Comerpickles:

978?

563blackghost:

I did not receive that. \(\Large\bf{\frac{89000000}{8700} \approx 1023}\)

Comerpickles:

Wait, where did all those zero's come from?

563blackghost:

it says `8.9 million`

Comerpickles:

Ohhhhh, forgot. Thank you. Could you help with 2 more please? If not thats totally alright I'd hate to bother

563blackghost:

Wait

563blackghost:

I added one extra zero. \(\large\bf{\frac{8900000}{8700}=102}\) Sorry about dat >.<

Comerpickles:

Sure, take all the time you need :)

Comerpickles:

Its alright, your answer was correct XD

563blackghost:

oh okie

563blackghost:

2 more? sure.

Comerpickles:

Are you sure? I'd hate to bother if you're busy

563blackghost:

its okie i can help with 2 more >.<

Comerpickles:

Okie dokie, bringing them up

Comerpickles:

Wait, wrong one

Comerpickles:

On second thought nevermind. I feel so bad asking because I know how busy you can be